### v1

# Intelligent Framework for Human-Robot Collaboration: Safety, Dynamic Ergonomics, and Adaptive Decision-Making

Abstract
The integration of collaborative robots into industrial environments has improved productivity, but has also highlighted significant challenges related to operator safety and ergonomics. This paper proposes an innovative framework that integrates advanced visual perception technologies, real-time ergonomic monitoring, and Behaviour Tree (BT)-based adaptive decision-making. Unlike traditional methods, which often operate in isolation or statically, our approach combines deep learning models (YOLO11 and SlowOnly), advanced tracking (Unscented Kalman Filter) and dynamic ergonomic assessments (OWAS), offering a modular, scalable and adaptive system. Experimental results show that the framework outperforms previous methods in several aspects: accuracy in detecting postures and actions, adaptivity in managing human-robot interactions, and ability to reduce ergonomic risk through timely robotic interventions. In particular, the visual perception module showed superiority over YOLOv9 and YOLOv8, while real-time ergonomic monitoring eliminated the limitations of static analysis. Adaptive role management, made possible by the Behaviour Tree, provided greater responsiveness than rule-based systems, making the framework suitable for complex industrial scenarios. Our system demonstrated a 92.5% accuracy in grasping intention recognition and successfully classified ergonomic risks with real-time responsiveness (average latency of 0.57 seconds), enabling timely robotic
Index Terms:
Human-robot collaboration, real-time ergonomics, adaptive decision-making, visual perception, Behaviour Trees, YOLO.
Model Params (M) FLOP (G) mAPval 50-95 (%) Latency (ms) Segmentation
YOLOv8-N 3.2 8.7 37.3 6.16 Yes (YOLO-seg)
YOLOv8-S 11.2 28.6 44.9 7.07 Yes (YOLO-seg)
YOLOv8-M 25.9 78.9 50.6 9.50 Yes (YOLO-seg)
YOLOv8-L 43.7 165.2 52.9 12.39 Yes (YOLO-seg)
YOLOv8-X 68.2 257.8 53.9 16.86 Yes (YOLO-seg)
YOLOv9t 2.0 7.7 38.3 N/A Yes (YOLO-seg)
YOLOv9s 7.2 26.7 46.8 N/A Yes (YOLO-seg)
YOLOv9m 20.1 76.8 51.4 N/A Yes (YOLO-seg)
YOLOv9c 25.5 102.8 53.0 N/A Yes (YOLO-seg)
YOLOv9e 58.1 192.5 55.6 N/A Yes (YOLO-seg)
YOLOv10-N 2.3 6.7 39.5 1.84 No
YOLOv10-S 7.2 21.6 46.8 2.49 No
YOLOv10-M 15.4 59.1 51.3 4.74 No
YOLOv10-L 24.4 120.3 53.4 7.28 No
YOLOv10-X 29.5 160.4 54.4 10.70 No
YOLO11-N 2.6 6.5 39.5 1.50 Yes (YOLO-seg)
YOLO11-S 9.4 21.5 47.0 2.50 Yes (YOLO-seg)
YOLO11-M 20.1 68.0 51.5 4.70 Yes (YOLO-seg)
YOLO11-L 25.3 86.9 53.4 6.20 Yes (YOLO-seg)
YOLO11-X 56.9 194.9 54.7 11.30 Yes (YOLO-seg)
EfficientDet (D3) 21.5 46.0 45.4 3.82 No
Faster R-CNN 41.3 87.5 49.8 6.38 No
Mask R-CNN 44.5 170.0 52.0 12.8 Yes
TABLE I: Comparison of COCO Dataset Architects. The performance of the examined models is compiled in the table, which further emphasizes how YOLO11 uniquely combines accuracy, speed, and segmentation capacity. This combination presents it as the ideal approach to handle the operational difficulties of collaborative environments, therefore enhancing both efficiency and safety.
I Introduction
Industrial automation has revolutionized manufacturing processes and enabled the integration of collaborative robotic systems, allowing robots and human operators to work together and enhancing the productivity and adaptability of production lines. This paradigm, termed Human-Robot Collaboration (HRC), has shown considerable productivity enhancements while also presenting intricate issues concerning safety, ergonomics, and flexibility in dinamics operational situations [1].
Figure 1: Schematic depiction of the proposed structure, illustrating the main modules and their interaction dynamics.
The prevention of Work-Related Musculoskeletal Disorders (WMSD) [2], which can arise due to incorrect postures or repetitive strain, is one of the most relevant issues. These disorders not only compromise the health of operators, but also negatively affect company productivity, increasing operating costs. As demonstrated by extensive studies on muscle fatigue in work environments [3], these issues require comprehensive monitoring solutions that can adapt to dynamic working conditions and provide timely interventions. Traditional technologies such as OWAS (Ovako Working Posture Analysis System) and REBA (Rapid Entire Body Assessment) offer useful tools for identifying ergonomic risks, but are inadequate for the continuous and dynamic monitoring required in today’s complex industrial environments.
To the best of our knowledge, no existing framework in the literature synergistically integrates advanced technologies such as visual detection, real-time ergonomic monitoring, and adaptive decision-making through Behavior Trees (BTs) for collaborative industrial settings. Current approaches address these challenges in isolation or only partially, limiting their practical applicability [4]. This work aims to bridge this gap by proposing an innovative framework that combines state-of-the-art technologies to enhance safety, ergonomics, and efficiency in industrial environments.
The proposed framework, illustrated in Fig. 1, stands out for its non-invasive nature and its ability to adapt to complex and dynamic operational scenarios. It integrates advanced visual detection technologies (YOLOv11 and OpenPose) for posture recognition, a modular decision-making system based on Behavior Trees (BTs) for dynamic human-robot role adaptation, and continuous ergonomic assessment methods to prevent physical risk situations. This integration approach addresses the safety challenges highlighted in industrial collaboration studies [1], while providing the adaptability required for varied manufacturing tasks and the non-invasiveness that traditional sensor-based systems [5] often lack. This unified approach enables real-time monitoring and optimization of human-robot interactions, enhancing both safety and overall productivity.
The remainder of this paper is organized as follows: Section II reviews the key related works to highlight existing gaps and motivate our approach. Section III provides a detailed description of the proposed framework, with a focus on its technological modules and system architecture. Section IV presents the experimental results, while Section V discusses potential future developments of the system.
II Related Works
In recent years, the field of human-robot collaboration has seen significant developments in several areas, such as visual perception, ergonomic evaluation and decision-making models to ensure safe and efficient robotic intervention. As highlighted by Villani et al. [4], these advancements have created new possibilities for intuitive and safe human-robot interactions in industrial settings, yet the integration of these technologies remains a significant challenge. Despite these advances, current solutions often suffer from limited integration of available technologies or take a static approach to ergonomics and safety. The need for more dynamic and adaptative approaches has been emphasized in multiple studies on work-related musculoskeletal disorders [3, 2], which identify real-time monitoring and adaptation as key factors in preventing occupational injuries.
II-A Visual Perception and Object Detection
The visual detection and segmentation of objects are crucial to ensure safety and efficiency in human-robot interactions [6]. In our framework, we chose to adopt YOLO11 [7], a state-of-the-art model that offers an excellent balance between accuracy, speed of inference and native segmentation capabilities, as reported in Table I.
Compared to its predecessors, YOLO11 [7] stands out due to significant optimisations in speed and accuracy. Models such as YOLOv8 [8], YOLOv9 [9] and YOLOv10 [10], while representing advances over previous generations, show limitations in handling complex scenarios or operational speed. At the same time, EfficientDet [11], although appreciated for its scalability and energy efficiency, is less competitive in dynamic contexts due to higher inference times and lack of native segmentation.
Even when compared to more complex models such as Mask R-CNN [12] and Faster R-CNN [13, 14], YOLO11 [7] stands out due to its lighter computational load, which makes it ideal for real-time applications where speed is essential to ensure effective responses. These features make it particularly suitable for complex collaborative scenarios, where rapid movements and frequent overlaps require highly responsive and accurate technology.
Integrated into our framework, YOLO11[7] demonstrated a remarkable ability to detect postures and objects even in scenarios characterised by complex movements, frequent overlaps and variable geometries. Its accuracy and speed proved essential to improve the tracking and monitoring of interactions, allowing the system to quickly adapt to operational requirements.
In conclusion, as is clear from Table I, YOLO11[7] represents a significant evolution from the original YOLO concept [6] and a significant step ahead of the alternatives due to its ability to combine high accuracy, accurate segmentation and operational speed. This evolution addresses key limitations identified in early object detection systems while maintaining the real-time performance critical for human-robot collaborative environments [1], where even milliseconds of delay can impact safety and operational efficiency. This combination makes it an ideal tool to address the challenges of human-robot collaboration, improving both safety and operational efficiency.
II-B Real-Time Tracking
Real-time tracking is crucial to ensure safety and fluidity in human-robot interactions. The Speed and Separation Monitoring (SSM) method proposed by Marvel et al. [15] monitors the speed and distance between human and robot to prevent collisions, but the lack of integration with advanced visual perception technologies limits its fluidity in complex environments . In contrast, Simple Online and Realtime Tracking (SORT) [16] offers a more responsive solution, using Kalman filtering [17] for rapid detection coupling, but suffers from reduced performance in the presence of occlusions or non-linear movements .
In our framework, the integration of YOLO11 with the Unscented Kalman Filter (UKF) [18] improves the handling of non-linear dynamics, overcoming the limitations of previous approaches such as SORT. This approach builds upon fundamental filtering principles [17] while extending their applicability to the complex, non-linear movements typical of human operators in industrial settings. This integration offers more reliable tracking in complex and rapidly changing industrial environments, , making operations safer and more efficient as emphasized by safety guidelines in collaborative robotic systems [1].
II-C Pose Detection and Action Recognition
Human pose and action recognition is essential for improving safety and efficiency in human-robot collaborations. While OpenPose [19] offers powerful pose estimation capabilities, its integration with ergonomic analysis systems remains challenging, as noted in human-robot collaboration surveys [4]. Although OpenPose [19] is widely used for real-time human joint detection, it does not offer dynamic ergonomic evaluation or decision-making based on classified actions.
In our previous work [20], we demonstrated that the SlowOnly network offers better performance in recognising slow, repetitive movements common in industrial settings than models such as SlowFast [21] and I3D [22] .
Lasota et al. [23], in an extensive survey of safety methodologies in human-robot interactions, highlight the importance of human action recognition for collision avoidance. However, most approaches are based on two-dimensional models, which do not adequately address the temporal and spatial complexity of human actions. Our framework, which integrates OpenPose 3D and SlowOnly, offers a more robust and efficient three-dimensional analysis to improve safety and ergonomics.
Cherubini et al. [24] explore collision avoidance in production scenarios, but do not integrate a three-dimensional system for action recognition, nor neural networks optimised for temporal recognition. In contrast, our approach exploits SlowOnly to accurately recognise human actions, improving the management of physical interactions in real time. This attention to temporal patterns in human movement aligns with the safety priorities outlined in collaborative manufacturing studies [1, 4], which identify action prediction as a key component of proactive safety systems. Peternel et al. [25] propose a system for the management of muscle fatigue during human-robot collaborations, but do not exploit networks optimised for the detection of complex and repetitive movements. Our framework fills this gap, improving ergonomic adaptation and robotic response through more sophisticated action recognition capabilities that can detect subtle movement patterns associated with fatigue and ergonomic risk [3].
II-D Dynamic Role Allocation with Behaviour Trees
Dynamic role allocation is crucial for efficient collaboration, especially in industrial settings where conditions can change rapidly. Rozo et al. [26] have employed probabilistic models to adapt roles based on human demonstrations, but these models struggle to handle dynamic industrial environments.
In our framework, we use Behaviour Trees (BT) [27], which offer a more modular and flexible decision-making structure, adapting robotic tasks to operational conditions detected in real time. Recent advances in unified architectures for dynamic role allocation and collaborative task planning in mixed human-robot teams [28] further demonstrate the importance of adaptable frameworks. Merlo et al. [29] proposed a similar framework based on ergonomic indicators, aimed at reducing the risk of musculoskeletal fatigue through intelligent role allocation. Our work extends this approach by integrating continuous monitoring of operators’ physical conditions to improve overall safety, addressing key concerns highlighted in comprehensive surveys of HRC methodologies [23] and providing a more responsive system for dynamic task environments.
II-E Real-Time Ergonomic Assessment
Traditional ergonomic assessment techniques, such as OWAS [30], RULA [31], REBA [32] and the NIOSH Lifting Equation [33], are based on manual observations and post-hoc analyses, and are unsuitable for continuous and dynamic monitoring of modern industrial environments. These methods have been systematically compared for their effectiveness in identifying potential work-related musculoskeletal disorders [34]. More recent studies, such as that of Ferraguti et al. [35] have proposed a solution to automate ergonomic assessment in HRC collaborations, but such approaches do not always succeed in continuously monitoring the physical condition of operators.
Our framework overcomes these limitations by integrating dynamic ergonomic analysis with advanced computer vision technologies such as OpenPose, enabling continuous monitoring of postures in real time [36]. This real-time analysis capability addresses a fundamental gap identified in traditional ergonomic evaluation methods [30, 33], which typically require manual observation and cannot adapt to rapidly changing work conditions. This approach not only prevents injuries related to incorrect posturesas cataloged in traditional ergonomic assessment methods [31, 32], but also enables an immediate adaptive response, improving safety and reducing operator muscle fatigue through interventions that align with established ergonomic principles [34].
Figure 2: Overview of the system’s perception and decision-making pipeline, showing the interplay between human pose tracking, action recognition, object detection, and ergonomic fatigue assessment.
II-F Constraints of Wearable Sensor Methodologies and Benefits of Computer Vision
Several studies have explored the use of wearable sensors to monitor ergonomic risk and assess operators’ movements in work contexts. For example, Santopaolo et al. [37] used inertial sensors and machine learning to classify biomechanical risks related to lifting, while Donisi et al. [38] combined wearable sensors with the NIOSH Lifting Equation to provide a detailed assessment of ergonomic risks in lifting tasks. Conforti et al. [5] have developed a system based on wearable sensors to monitor operators’ movements, demonstrating the effectiveness of these systems for collecting detailed posture and movement data.
Despite their accuracy, wearable sensor-based approaches have significant limitations. Sensors may be invasive, interfering with operators’ movements and requiring ongoing management for charging, calibration, and maintenance. Moreover, such systems can increase operational costs, especially in large-scale industrial environments where every worker must be equipped with physical devices.
In contrast, our framework based on artificial vision offers a non-invasive solution for continuous ergonomic assessment. Using technologies such as YOLO11[7] for object detection and OpenPose for human posture analysis, the system monitors operators’ posture and movements in real-time without requiring the use of physical devices. This approach builds upon established object detection principles [6] while overcoming the limitations of traditional monitoring systems, providing a scalable solution that can be deployed across multiple workstations without additional hardware costs per operator. This allows a more natural assessment of ergonomic conditions, dynamically adapting to changes in operators’ movements.
In addition, computer vision allows for scalable coverage in complex environments, monitoring multiple operators and robots simultaneously without the need for additional sensors. The system can identify incorrect postures or risky movements and intervene in real-time, reducing the risk of repetitive motion or incorrect posture-related injuries, as confirmed by previous studies on musculoskeletal disorders [3, 33].
In summary, although wearable sensors offer high accuracy, our computer vision-based approach has significant advantages in terms of flexibility, non-invasiveness and scalability, making it particularly suitable for dynamic industrial settings where real-time ergonomic assessment is required. This approach aligns with the evolution of industrial safety paradigms [4, 1] toward more integrated and adaptable solutions that can accommodate the full range of human-robot collaborative scenarios while minimizing disruption to existing workflows.
II-G Proposed Framework Improvements
Compared to previous works, particularly those that address individual aspects of human-robot collaboration [15, 19, 27] rather than providing an integrated solution, our framework has the following advantages:
  • Dynamic Ergonomic Evaluation: Provides continuous and real-time evaluation, improving injury prevention due to incorrect postures or repetitive movements. This continuous monitoring represents a significant advancement over the static approaches criticized in previous studies [35] and addresses key concerns related to musculoskeletal disorders [3], with our experimental results showing effective classification of ergonomic risks in real-time (average latency of 0.57 seconds).
  • Advanced Action Perception: The integration of OpenPose with SlowOnly allows for an in-depth understanding of human actions, identifying movements that could lead to excessive physical stress. This integration addresses the challenges of temporal recognition identified in prior work [24] and extends the capabilities of standard pose estimation systems [19], achieving 95.83% accuracy in action recognition during our validation tests.
  • Intelligent Automation: The behaviour tree-based decision-making system allows selective and adaptive intervention by the robot, optimising operational efficiency without interfering with human activities. Our behavior tree implementation extends existing role allocation frameworks [29] with enhanced adaptability, while addressing safety concerns highlighted in comprehensive HRC surveys [23, 4], demonstrating rapid response times (0.07 seconds on average) to critical ergonomic conditions.
III Proposed Framework Architecture
The proposed framework, in Fig. 2, combines advanced computer vision techniques, human action recognition, and ergonomic evaluation, enabling synergy between human operator and robot in collaborative environments. This architecture enables real-time monitoring of operators’ operating conditions, with an adaptive decision-making system managing dynamic task allocation between human and robot via a Behaviour Tree (BT).
The modular design of our framework allows for component-level improvements and adaptation to various industrial contexts without requiring complete system redesign, addressing a key limitation of monolithic approaches identified in previous research [4]. Integration between different technologies, such as computer vision and ergonomic evaluation systems, is essential to improving operations safety and efficiency. In the following, we analyze each key module, explaining how they interact with each other and contribute to the overall operation of the framework.
III-A Visual Perception Module
The Visual Perception Module is responsible for detecting, tracking, and calculating the physical characteristics of objects and operators’ hands. Data flow from YOLO11 and tracking via Unscented Kalman Filter (UKF) provides parallel inputs that feed into the decision-making system and ergonomic evaluation.
III-B YOLO11: Object Detection
For object detection and the generation of accurate segmentation masks, the YOLO11 model was used, which provides detailed data on the position and size of detected objects, which is essential for ensuring safety in human-robot collaboration environments. The output of the model is represented by a vector $\mathbf{y}$, which includes position, size, confidence and segmentation of objects:
$\mathbf{y}=(x,y,w,h,c,\mathbf{p},\mathbf{m})$
where $(x,y)$ represents the coordinates of the center of the bounding box, $w$ and $h$ are the width and height of the bounding box, $c$ denotes the confidence of the object’s presence, $\mathbf{p}$ is the vector of classification probabilities, and $\mathbf{m}$ is the binary segmentation mask.
The model is optimized through a loss function that balances location, confidence, classification, and segmentation:
$\mathcal{L}_{\text{totale}}=\mathcal{L}_{\text{box}}+\lambda_{\text{conf}}\mathcal{L}_{\text{conf}}+\lambda_{\text{cls}}\mathcal{L}_{\text{cls}}+\lambda_{\text{seg}}\mathcal{L}_{\text{seg}}$
Where $\lambda_{text{conf}}$, $\lambda_{text{cls}}$ and $\lambda_{text{seg}}$ are scalar weights that balance the relative importance of each component of the loss.
This formulation allows the system to detect and segment objects in complex scenarios, ensuring accurate separation even in the presence of overlaps or irregular shapes.
III-C Calculation of Area, Volume, and Volumetric Weight: Input from YOLO11
Using the data provided by YOLO11, the system performs the parallel calculation of the area, volume, and volumetric weight of objects, which are key inputs for ergonomic evaluation using OWAS.
Area:
The area of the object in the 2D projection is calculated as:
$A_{\text{object}}=w\cdot h$
where $w$ and $h$ are the width and height of the bounding box.
Volume:
The volume of the object is estimated by multiplying the area by the average depth $D_{\text{object}}$detected in the bounding box:
$V_{\text{object}}=A_{\text{object}}\cdot(D_{\text{max}}-D_{\text{object}})$
where $D_{\text{max}}$ represents the maximum distance and $D_{\text{object}}$ is the average depth detected by the camera RealSense.
Volumetric Weight:
Once the volume is estimated, the volumetric weight is calculated by dividing the volume by a standard conversion factor:
$P_{\text{volumetric}}=\frac{V_{\text{object}}}{\text{conversion factor}}$
where conversion factor is a standard value of 6000 used in logistics companies to calculate volumetric weight.
This input is critical to the calculation of the OWAS score, where the weight of objects handled by operators is one of the key factors in assessing ergonomic risk.
Figure 3: OWAS evaluation and result interpretation [39].
III-C1 Kalman Unscented Filter (UKF)
In parallel with calculating the physical properties of objects, the
III-C2 Unscented Kalman Filter
(UKF) is used to track the positions and movements of objects and operators’ hands in real time. UKF is particularly well suited to handle nonlinear dynamics typical of complex movements in industrial environments.
Tracking is handled through a state vector $\mathbf{x}$ that includes the positions, velocities, and orientations of objects and hands:
$\mathbf{x}=\begin{bmatrix}\mathbf{p}_{\text{parcel}}\\ \mathbf{v}_{\text{parcel}}\\ \mathbf{q}_{\text{parcel}}\\ \mathbf{p}_{\text{hand\_left}}\\ \mathbf{v}_{\text{hand\_left}}\\ \mathbf{q}_{\text{hand\_left}}\\ \mathbf{p}_{\text{hand\_right}}\\ \mathbf{v}_{\text{hand\_right}}\\ \mathbf{q}_{\text{hand\_right}}\end{bmatrix}$
Here, $\mathbf{p}$ denotes the position in 3D space, $\mathbf{v}$ the linear velocity in 3D space, and $\mathbf{q}$ the orientation expressed in quaternions to avoid the ambiguities that arise in the use of Euler angles.
The state update is done by a transition function $\mathbf{f(x)}$ that predicts the evolution of the variables over time. This function takes into account the linear and rotational dynamics of the system and is expressed as follows:
$\mathbf{f(x)}=\begin{bmatrix}\mathbf{p}_{\text{parcel}}+\mathbf{v}_{\text{parcel}}\cdot dt\\ \mathbf{v}_{\text{parcel}}\\ \mathbf{q}_{\text{parcel}}+\frac{1}{2}\mathbf{q}_{\text{parcel}}\otimes\mathbf{\Omega}(\boldsymbol{\omega}_{\text{parcel}})\cdot dt\\ \mathbf{p}_{\text{hand\_left}}+\mathbf{v}_{\text{hand\_left}}\cdot dt\\ \mathbf{v}_{\text{hand\_left}}\\ \mathbf{q}_{\text{hand\_left}}+\frac{1}{2}\mathbf{q}_{\text{hand\_left}}\otimes\mathbf{\Omega}(\boldsymbol{\omega}_{\text{hand\_left}})\cdot dt\\ \mathbf{p}_{\text{hand\_right}}+\mathbf{v}_{\text{hand\_right}}\cdot dt\\ \mathbf{v}_{\text{hand\_right}}\\ \mathbf{q}_{\text{hand\_right}}+\frac{1}{2}\mathbf{q}_{\text{hand\_right}}\otimes\mathbf{\Omega}(\boldsymbol{\omega}_{\text{hand\_right}})\cdot dt\end{bmatrix}$
In this formulation:
  • $\mathbf{p}$ is the updated position as a function of linear velocity $\mathbf{v}$ and time $dt$,
  • $\mathbf{v}$ represents velocity, held constant within the update interval $dt$. This choice is justified by the high update rate of the system, which makes speed changes between two consecutive updates negligible,
  • $\mathbf{q}$ represents the updated orientation using the quaternion product with the angular velocity $\boldsymbol{\omega}$ and time $dt$.
The predicted estimate of the state, after updating, is given by the transition function:
$\hat{\mathbf{x}}_{k+1}=f(\mathbf{x}_{k})$
State correction
After estimating the future state with the transition function, the UKF compares this estimate with actual observations obtained from depth sensors, such as the RealSense D435 camera. The measurement function $\mathbf{h(x)}$maps the estimated state onto observables, i.e. the position and orientation of objects and hands:
$\mathbf{h(x)}=\begin{bmatrix}\mathbf{p}_{\text{parcel}}\\ \mathbf{q}_{\text{parcel}}\\ \mathbf{p}_{\text{hand\_left}}\\ \mathbf{q}_{\text{hand\_left}}\\ \mathbf{p}_{\text{hand\_right}}\\ \mathbf{q}_{\text{hand\_right}}\end{bmatrix}$
The measurement predicted at time $k$,
$\hat{\mathbf{z}}_{k}$
is obtained by applying the measurement function to the state predicted at time $k$:
$\hat{\mathbf{z}}_{k}=\mathbf{h}(\hat{\mathbf{x}}_{k})$
Subsequently, the new actual observations ${\mathbf{z}}_{k}$ are compared with those predicted estimates. The error between the predicted estimate $\hat{\mathbf{z}}_{k}$ and the actual observations $\mathbf{z}_{k}$ is used to update and correct the estimated state, via the Kalman gain $K$.The final correction of the state $\mathbf{x}_{k+1}$, is made with the following equation:$k$:
$\mathbf{x}_{k+1}=\hat{\mathbf{x}}_{k+1}+K(\mathbf{z}_{k}-\hat{\mathbf{z}}_{k})$
In this way, the estimated state is updated using new sensory observations, improving accuracy in real time.
Covariance Matrices
The stability and accuracy of the system strongly depend on a proper definition of the covariance matrices. The process noise covariance matrix $\mathbf{Q}$ represents the uncertainty in the system model, while the observation noise matrix $\mathbf{R}$ handles the uncertainty associated with sensor measurements:
$\mathbf{Q}=\text{diag}([100,\dots,100])\quad\text{e}\quad\mathbf{R}=\text{diag}([0.01,\dots,0.01])$
The initial state covariance matrix $\mathbf{P}$, defined as a scaled identity matrix, also plays a key role in the accuracy of the first estimates:
$\mathbf{P}=0.01\cdot\mathbf{I}_{30}$
The correct choice of these parameters is essential to ensure that the UKF system maintains stability and accuracy even in the presence of uncertainties in the process or measurements, minimizing estimation error through the entire operational cycle.
III-D Human Actions Recognition Module
The
III-E Human Actions Recognition Module
integrates
III-F OpenPose
for pose detection and
III-G SlowOnly
for large time-scale action recognition.
III-H OpenPose: Pose Detection
OpenPose detects the joints of the human body in 3D using a perspective projectionciteiodice2022learning. Each i-th joint is represented by the coordinates $(x,y,z)$ and a confidence score $c$:
$\mathbf{J}=\left\{(x_{i},y_{i},z_{i},c_{i})\right\},\,i=1,2,\dots,N$
The collected data are subsequently analyzed by the SlowOnly module.
III-I SlowOnly: Temporal Analysis
The SlowOnly model, a three-dimensional convolutional network (3D ConvNet), analyzes the slow, repetitive movements of operators. The input is a sequence of video frames $F_{text{in}}$, sampled at regular intervals:
$F_{\text{in}}=\left[I_{t},I_{t+\Delta t},I_{t+2\Delta t},\dots,I_{t+T\Delta t}\right]$
where $I_{t}$ is the video frame at time $t$, $\Delta t$ the temporal sampling interval, and $T$ the total number of frames in the sequence.
Three-dimensional convolution applied to this sequence extracts the temporal and spatial features:
$F_{\text{out}}=\text{Conv3D}(F_{\text{in}},K,T)$
III-J OWAS-Based Ergonomic Assessment
The Ovako Working Posture Analysis System (OWAS), classifies operators’ postures into risk categories, allowing the system to monitor and intervene when postures are detected that could lead to musculoskeletal injuries.
III-K Classification of Postures
The classification of operator postures, in Fig.3, involves four categories:
  1. 1.
    Posture of the back: straight, flexed or rotated.
  2. 2.
    Posture of the arms: above or below shoulder level.
  3. 3.
    Position of legs: standing, sitting, kneeling or squatting.
  4. 4.
    Weight lifted: amount of load lifted by the operator.
Each combination of these categories is ranked according to an ergonomic risk score, which indicates the level of urgency for corrective action:
  • Level 1: No ergonomic risk, no need for intervention.
  • Level 2: Moderate risk, long-term corrective interventions may be needed.
  • Level 3: High risk, short-term corrective intervention required.
  • Level 4: Extremely risky, immediate corrective action is required.
III-L Calculation of OWAS Score
The overall OWAS score for a posture is calculated as a weighted combination of the scores of the different categories (back, arms, legs, and load). The formula for calculating the overall score is:
$P_{\text{owas}}=\alpha\cdot P_{\text{back}}+\beta\cdot P_{\text{arm}}+\gamma\cdot P_{\text{leg}}+\delta\cdot P_{\text{weight}}$
where:
  • $P_{\text{back}}$, $P_{\text{arm}}$, $P_{\text{leg}}$, $P_{\text{weight}}$ are the scores assigned to each category,
  • $\alpha$, $\beta$, $\gamma$, $\delta$ are the weights assigned according to the severity of risk for each category.
Figure 4: Behaviour Tree diagram for dynamic role allocation between robot and operator based on physical and operational conditions..
III-M Integration of OWAS into the Framework
OWAS is integrated directly into the framework to monitor operators’ posture in real time, using data from OpenPose and SlowOnly. Data from body joints, collected by OpenPose, are mapped to OWAS categories (back, arms, legs, and load) to automatically assess operator posture.
The assessment process consists of the following steps:
  1. 1.
    Postural data collection: OpenPose detects the position of the body joints and builds a virtual skeleton.
  2. 2.
    Posture Classification: The collected information is mapped to OWAS categories.
  3. 3.
    Calculation of OWAS score: The overall OWAS score is calculated for each combination of detected postures and loads.
  4. 4.
    Dynamic Intervention: If the OWAS score exceeds a predefined threshold, the system suggests the operator to change the posture or stop the activity to prevent injury.
III-N Behaviour Tree (BT)
based automation. Within the proposed framework, dynamic role allocation between human operator and robot is managed through a Behaviour Tree (BT). This decision structure allows the system to adapt robotic behavior based on operational conditions, operator actions, and real-time sensory inputs. Unlike traditional approaches, such as finite-state diagrams or Markovian decision models, Behaviour Tree offers greater flexibility and modularity, which is particularly relevant in dynamic industrial settings where human-robot collaboration requires fast and accurate decisions.
III-O Behaviour Tree structure
The Behaviour Tree implemented in the framework is hierarchically structured and composed of different types of nodes:
  • Action Nodes: They perform concrete operations, such as “grabbing an object” or “moving to a destination.” These nodes allow the robot to perform the required actions based on the detected conditions.
  • Conditional Nodes: They evaluate whether or not the action can proceed by monitoring parameters such as operator biomechanical fatigue or completion of a sequence of human movements. For example, a conditional node might check whether the lifted package has an acceptable weight for handling.
  • Composite Nodes: These nodes combine multiple actions and conditions, creating complex logic flows. Composite nodes used include:
    • Sequence (Sequence): Nodes are executed in sequential order until all actions are completed correctly. If a node fails, execution stops.
    • Fallback: Nodes are executed sequentially until one succeeds. This node type is particularly useful for handling alternative situations where the primary action is unavailable or fails.
Figure 4 illustrates the structure of the decision tree used in the framework to dynamically allocate roles between the operator and the robot based on operational conditions and the ergonomic situation.
This modular structure allows the system to perform complex actions and make dynamic decisions based on conditions sensed in real time. For example, if the human operator is fatigued or has not completed a certain sequence of movements, the BT automatically triggers alternatives such as robot intervention to lift and move the package.In this way, the BT manages role allocation flexibly, maintaining operator safety and operational efficiency.
III-P Execution of the Behaviour Tree
The execution of the Behaviour Tree starts from the root node, which manages the operational flow of the robot. The root sequence starts a series of conditional and action nodes, including:
  • GoHoming: Verifies that the robot is in the correct starting position before starting any action.
  • CheckHumanSequenceCompleted: Evaluates whether the operator has completed the intended sequence of movements.
  • CheckBiomechanicsFatigue: Monitors operator biomechanics fatigue based on ergonomic data collected in real time.
If all conditions are met, the BT proceeds with a series of specific actions, such as:
  • MoveToParcel: The robot moves toward the parcel to be lifted.
  • GraspParcel: It performs the action of grabbing the parcel.
  • MoveToContainer: Moves the parcel to the designated container.
  • ReleaseParcel: Releases the parcel into the container.
In case the ergonomic condition of the operator is not optimal, such as due to muscle fatigue or poor posture, the BT triggers corrective actions, such as suggesting the operator to stop the activity or take a secondary role. This ensures safe and efficient interaction between human and robot.
III-Q Advantages of using Behaviour Tree in the framework
The adoption of the Behaviour Tree in the implemented framework offers several advantages over other control techniques, including:
  • Modularity: Each node in the BT is independent and modifiable, allowing behaviors to be added or removed without affecting the entire system.
  • Flexibility: The system can dynamically adapt to changes in the operator’s physical condition or variations in operational activities, responding in real time to factors such as musculoskeletal fatigue or variation in load weight.
  • Real-Time Responsiveness: Through integration with visual perception modules (e.g., YOLO and OpenPose) and ergonomic evaluation, the BT enables the robot to make quick and accurate decisions, optimizing both operator safety and operational efficiency.
IV Experiments and Results
To evaluate the overall effectiveness of the proposed framework in human-robot collaboration, experiments were designed and conducted in a controlled laboratory environment and in simulation. The laboratory experiments replicated realistic operational scenarios, focusing on ergonomic monitoring, action detection, visual perception, and overall integration of the framework components. In parallel, the Behaviour Tree (BT)-based decision-making module was tested in simulation to analyze its dynamic and adaptive management capability in a wide range of complex scenarios that are difficult to replicate in the laboratory.
IV-A Experiment: Grasping Action Detection and Volumetric Weight Estimation with Variable Size Parcels
TABLE II: Global performance in terms of mAP
. Model mAP@50 (%) mAP@50:95 (%) YOLOv8x-seg 77.6 71.2 YOLOv9-seg 77.7 72.4 YOLO11x-seg 77.8 72.4
TABLE III: Performance per class: accuracy, recall, and mAP
Class Metrics YOLOv8x-seg (%) YOLOv9-seg (%) YOLO11x-seg (%)
Parcel Precision 88.0 85.6 89.8
Recall 79.3 80.9 81.1
mAP@50 89.4 89.4 89.3
mAP@50:95 86.7 86.0 86.8
Hand Precision 89.8 89.6 89.8
Recall 59.9 60.9 61.1
mAP@50 65.7 66.1 66.3
mAP@50:95 57.3 53.2 58.0
In this preliminary phase of the framework, the main objective was the recognition of grasping intention and estimation of volumetric weight of parcels of varying sizes. The dataset used included 15,160 images annotated with the Segment Anything Model (SAM), ensuring accurate segmentation of hands and parcels into diverse configurations.Images were divided into training (80%), validation (10%) and testing (10%), ensuring a balanced distribution for robust evaluation.
Three segmentation models were analyzed: YOLOv8x-seg, YOLOv9-seg e YOLO11x-seg, trained with the same configuration parameters and evaluated on two main classes, , ”parcel” and ”hand”. As shown in Table II, YOLOv8x-segachieved a mAP@50 of 77.6% and a mAP@50:95 of 71.2%, demonstrating a good balance between precision and recall. YOLOv9-seg slightly improved the mAP@50 (77.7%) and maintained the mAP@50:95 at 72.4%, with more efficient latency management. However, YOLO11x-seg outperformed both, achieving a mAP@50 of 77.8% and a mAP@50:95 of 72.4%, showing greater robustness in complex scenarios.
Table III analyzes performance by class. For the class ”parcel”, YOLOv9-seg and YOLO11x-seg produced similar results, with YOLO11x-seg slightly outperforming YOLOv9-seg in terms of recall and mAP@50:95 (86.8% versus 86.0%). For the class ”hand”, YOLO11x-seg stood out significantly, achieving a mAP@50:95 of 58.0%,compared with 53.2%for YOLOv9-seg and 57.3% for YOLOv8x-seg. This result highlights the ability of YOLO11x-seg to handle pose variability, occlusions, and structural complexity.
For grasping intention recognition, 40 trials were conducted on dynamic sequences, combining the segmentation model with the probabilistic Kalman Unscented (UKF) filter. The system correctly identified grasping in 37 cases out of 40, with an overall accuracy of 92.5%. For volumetric estimation, 4 parcels of varying sizes and shapes were tested, each in 10 trials for a total of 40 measurements. The estimated volumetric weights, obtained from the segmented masks and depth data provided by the RealSense D435 sensor, were compared with the actual weights. Percentage errors were calculated using the formula:
$E_{p}=\frac{|P_{v,\text{st}}-P_{v,\text{re}}|}{P_{v,\text{re}}}\times 100$
where $P_{v,\text{st}}$ e $P_{v,\text{re}}$ represent estimated and actual volumetric weights, respectively. The overall mean percent error over all trials was calculated as 17.58%, with greater variability observed in the extreme size parcels, attributable to the limitations of the sensor in terms of spatial resolution and surface reflectivity.
In conclusion, the model comparison confirmed that YOLO11x-seg represents the best performing solution for the operational context considered. Its combination of robustness, accuracy, and ability to handle difficult conditions, such as detection of moving or partially occluded hands, makes it ideal for grasping and volumetric estimation applications. These results highlight the advantages of integrating a deep learning model with probabilistic tracking algorithms, demonstrating superior robustness compared to the isolated use of neural networks. However, volumetric weight estimation showed sensitivity to the quality of depth data provided by the RealSense D435 sensor, with a reduction in accuracy due to limitations in spatial resolution or surface reflectivity. This suggests that further improvements could be achieved through refinement of sensor calibration or integration of data from multiple sensors to compensate for current limitations.
IV-B Experiment: Ergonomic Monitoring, Risk Classification and Robotic Intervention via Behavior Tree
The objective of this experiment is to evaluate the framework’s ability to monitor operator postures in real time, classify ergonomic risk, and manage task transfer to the robot in the presence of critical conditions. During testing, the operator performed movements typical of an industrial environment, such as bending and lifting, while the system monitored posture using OpenPose and classified ergonomic risk using the Ovako Working Posture Analysis System (OWAS) method.
When medium-high-risk (OWAS class 3) or high-risk (OWAS class 4) postures were detected, or volumetric weight in excess of allowable limits, the system generated a biomechanical fatigue message. This message was processed by a Behavior Tree (BT), which combined ergonomic data, action recognition results, and volumetric weight values to determine whether robotic intervention should be activated. This approach ensured modular and responsive decision making, allowing a smooth transition between the operator and the robot.
The data collected during the experiment showed that 12,8% of the postures belonged to OWAS class 1 (no risk), l’84,6% belonged to OWAS class 2 (moderate risk), no postures were classified in OWAS class 3 (medium-high risk), and 2,6% belonged to OWAS class 4 (high risk), as shown in Fig. 6. These results confirm the system’s ability to distinguish between ergonomically safe and hazardous postures, demonstrating the effectiveness of the framework in monitoring and classifying risk.
Temporally, the system maintained an average update rate of 14.99 Hz, with an average latency of 0.57 seconds for pose monitoring, as shown in Fig. 6. These values ensured that any postural changes were detected and analyzed in a timely manner. In addition, the average response time to risky situations was 0.07 seconds, as shown in Fig. 8, demonstrating the efficiency of the framework in identifying critical conditions and responding quickly.
An analysis of the postures over time windows of four samples, Fig. 8 revealed that the median of the postures remained in OWAS class 2 (moderate risk), with temporary peaks in OWAS class 4 during particularly heavy lifting or postures held for a long time. This underscores the system’s ability to accurately detect critical moments, providing useful decision support for operations management.
These results highlight not only the framework’s efficiency in detecting and analyzing critical conditions, but also its role in supporting a seamless operational transition. The reduced overall latency, as calculated through pose monitoring and response time to risky situations, ensured that the transfer of activities to the robot was smooth and responsive. This level of operational readiness, combined with the system’s ability to integrate ergonomic and actionable data, represents a significant step toward optimizing human-robot collaboration.
Figure 5: Distribution of postures classified according to OWAS classes during the experiment. Most of the postures belong to OWAS class 2 (moderate risk), with a small percentage in OWAS 4 (high risk).
Figure 6: Update latencies of monitored poses. An average latency of 0.57 seconds ensures timely detections and supports rapid interventions.
Figure 7: Average response times of the framework to critical conditions. The fast response (0.07 seconds on average) highlights the system’s efficiency in making real-time decisions.
Figure 8: Time trend of the median of the analyzed poses according to OWAS. Predominance of moderate risk (OWAS class 2) with peaks in OWAS class 4 during heavy lifting.
IV-C Experiment: Recognizing Actions with SlowOnly Model
Figure 9: Trend of classification loss during the 200 training epochs. The curve shows stable convergence, suggesting progressive learning of the model.
Figure 10: Increased top-1 accuracy on the validation set during the 200 training epochs. Accuracy shows continuous improvement, highlighting the effectiveness of model optimization..
This experiment evaluated the framework’s ability to recognize and classify actions performed by an operator in a simulated industrial environment by exploiting the SlowOnly model. The model, based on a ResNet3D network with SlowOnly architecture and 50-level depth, was pre-trained on the HRI30 dataset, specifically developed to recognize actions in human-robot interaction scenarios. During the experiment, the model was configured to classify three classes of actions: lifting, carrying, and repositioning packages.
The framework uses an innovative approach that combines action recognition with ergonomic and volumetric analysis, integrating the results within a Behavior Tree (BT) to determine possible robotic intervention. Video sequences were preprocessed to 8-frame clips, uniformly sampled with a time interval of four frames, and normalized using ImageNet mean values and standard deviations. Each clip was subjected to resizing, center cropping, and augmentation operations with random color changes and horizontal flips to improve the robustness of the model. Training was conducted on an NVIDIA RTX 3090 GPU, using a batch size of 16 videos per GPU.
The optimization process used the AdamW algorithm with an initial learning rate of 0.001, adjusted through a Cosine Annealing strategy with warmup for the first 1,000 steps. Training was performed for 200 epochs, with the cross-entropy loss function. During the validation phase, the model achieved an accuracy of 95,83%, as shown in Fig. 10, demonstrating an excellent ability to generalize over the test data.
The behavior of loss during training shows rapid initial convergence, with stabilization at 0.129 toward the later epochs, as shown in Fig. 10. This result reflects the effectiveness of the model setup and training pipeline in gradually reducing the classification error. The use of short 8-frame clips proved particularly effective in capturing action movements without introducing significant computational complexity, while integration with the BT ensured that the classified data were exploited for real-time robotic decisions.
This model and framework configuration was further validated by the previous evaluation on HRI30, where SlowOnly was shown to outperform alternative architectures such as I3D and TSM in action recognition in industrial settings. The model’s robustness to occlusions and its high accuracy make it an ideal choice for collaborative human-robot scenarios, where accurate and timely action recognition is crucial.
The results of this experiment confirm the effectiveness of the proposed approach, which combines action recognition, ergonomic analysis, and volumetric evaluation, integrating them into a modular and responsive decision-making framework. The framework showed a remarkable ability to adapt to operational variables, providing decisive support for safety and efficiency in collaborative industrial operations.
V Conclusions and Future Work
In this work, an innovative framework for human-robot collaboration was proposed, combining advanced visual sensing technologies, real-time ergonomic monitoring, and Behaviour Tree (BT)-based adaptive decision-making. Experimental results show that the system significantly improves operator safety and efficiency of operations, providing a scalable and modular solution for complex collaborative environments.
To our knowledge, the framework represents a novelty in the existing literature, synergistically integrating advanced perception technologies and real-time decision-making. Compared with traditional methods based on static rules or post-hoc ergonomic analysis, our approach introduces greater flexibility and adaptability, improving both the safety and well-being of operators. The noninvasive nature of the system, which eliminates the need for wearable physical sensors, also helps to reduce operational costs and improve the naturalness of human-robot interaction.
Despite the encouraging results, the work has some limitations. The experiments, conducted in a controlled laboratory and in simulation, need to be extended to real industrial scenarios to evaluate the robustness of the system under varying operating conditions. In particular, the Behaviour Tree module requires practical validation to demonstrate its effectiveness in dynamic human-robot role management. In addition, the absence of direct comparisons with established technologies presents an opportunity to further position the framework in the state of the art.
V-A Future Work
To address the limitations that have emerged and broaden the impact of the framework, future work will focus on:
  • Validation in Real Environments: Conducting field tests in operational industrial settings to demonstrate the robustness and scalability of the system on a large scale.
  • Comparisons with Existing Solutions: Conduct quantitative comparisons with established academic technologies using standard benchmarks to further validate the framework’s contribution.
  • End-User Engagement: Collaborate with industrial operators and supervisors to gather feedback on the usability, efficiency, and acceptability of the framework, improving its configuration based on specific user needs.
  • Expansion for Complex Scenarios: Extend the system to handle multi-operator and multi-robot interactions, improving collaborative efficiency in dynamic and high-variability environments.
  • Integration of Advanced Sensors: Experiment with next-generation sensors, such as high-resolution RGB-D cameras and multi-sensor configurations, to further improve the accuracy of visual monitoring and volumetric estimation.
In conclusion, the proposed framework represents a major breakthrough in human-robot collaboration, combining safety, efficiency and adaptability in a single integrated solution. Its modular and scalable architecture makes it particularly suitable for a wide range of industrial applications, including logistics, manufacturing, and healthcare. With validation in real-world environments and enrichment with additional functionality, the system has the potential to become a benchmark solution, improving the productivity and well-being of operators in modern manufacturing settings.